Calculate the force of gravity on a 20 kg mass at Earth's surface. The mass of Earth is 6. 1024 kg, and its radius is 6.4 106 m.?
1
Expert's answer
2012-10-24T10:36:14-0400
The magnitude of the gravity force acting on a 20 kg mass at Earth's surface is

F = G*m*M/r² =

= 6.67384×10^(-11)[m³/kg/s²] * 6×10^24[kg] * 20[kg] / (6.4*10^6[m])² =

& asymp; 195.5227[N].